# Rheology of dense vibrated granular flows: non-monotonic response controlled by granular temperature

### v1

@@ III A unified Picture
Now we point out that the observation of a non-monotonic behavior of $\eta$ as a function of the shaking frequency (Sec. II.1) is in striking contrast with the idea that the rheological response of the granular medium is mainly controlled by $A^{2}/p$. While the viscosity weakening branch can be understood by combining the acceleration threshold $\Gamma=1$ and the $A^{2}/p$ scaling, the raising branch $\eta(\Gamma\geq\Gamma_{\text{m}})$ is not. Indeed, simulations performed at $\Gamma\geq\Gamma_{\text{m}}$ (varied through $f$), with the same amplitude and pressure can manifest effective viscosities that vary over different orders of magnitude, see Fig. 3c. At the same time, we also verified that the kinetic energy of the system does not substantially depend on $p$ (not shown), thus the effect of $p$ on the effective viscosity (see Fig. 3a) can’t be explained with a dependence of $K$ on $p$. Based on our numerical analysis, we define an energy scale related to the confining pressure $K_{p}=p\pi r^{3}$ (i.e. the work necessary to move of a distance $r$ against a constant force $p\pi r^{2}$) and propose the energy ratio $\bar{K}=K/K_{p}$ as the main variable controlling the effective viscosity of the system. In Fig. 3d, we plot $\eta(\Gamma\geq\Gamma_{\text{m}})$ obtained for all the compatible frequencies, amplitudes and pressures varied in our analysis, showing a good collapse on the master curve $\eta\propto\bar{K}^{\alpha}$ with $\alpha\sim-2$.
@@
We point out that the ratio of energies defining $\bar{K}$ reminds of the competition between thermal agitation and activation energy barriers that characterizes the viscosity divergence in glass-forming supercooled liquids [debenedetti2001supercooled] with, however, the important difference that in these systems viscosity divergences are usually exponential or even steeper while in our case we observe a power law. This qualitative analogy with thermal systems is nevertheless conceptually useful to understand the relationship between rheological response and $\bar{K}(A,f,p)$ in our granular setup. A peculiarity of thermal systems is that temperature can be directly controlled by equilibrating the system with a thermal bath. Then, as a consequence of thermodynamic equilibrium, also the average kinetic energy of the system is fixed and proportional to the temperature. Instead, in a granular system, the average kinetic energy $K(A,f)$ depends non-trivially on the driving parameters. In other words, the relationship between injected and adsorbed energy is much more complex when thermodynamic equilibrium is violated, as in granular systems. Therefore, to understand how the driving parameters $A$ and $f$ control the rheology of a granular system it is important to first clarify how they act on its granular temperature $K$. Our numerical analysis shows that, since $\eta$ depends non-monotonously on $f$, there is no simple scaling in the form $\propto A^{\beta}f^{\lambda}$, with $\beta$ and $\lambda$ arbitrary exponents, that could predict the behavior of the effective viscosity. Panel c and d of Fig. 3 show in fact that $\eta$ depends on $A$ and $f$ through $K(A,f)$, which embodies the non-monotonic dependence on $f$.
@@ IV Conclusions and link with previous studies
In this paper we investigated the rheology of a dense vibrated granular system by means of realistic numerical DEM simulations of a stress imposed vane rheometry setup. We varied the vibration amplitude $A$ and frequency $f$ of the vertical sinusoidal shaking as well as the confining pressure $p$. We found that the variable controlling the effective viscosity is $\bar{K}=K/K_{p}$ where $K(A,f)$ is the typical agitation energy of a grain in the cage formed by the surrounding particles and $K_{p}\propto p$ is an energy scale set by the confining pressure that could be interpreted as the energy needed to escape from such a cage. Our proposed scaling takes into account the non-monotonic effect of $f$ on $\eta$ through the energy transfer mechanism discussed in Ref. [Plati2021Getting] and Appendix B.
@@
We now discuss the connection between our study and previous works on the rheology of vibrated granular materials in simplified geometries. We start referring to the numerical study reported in [Clark2023] which focuses on the macroscopic friction coefficient $\mu$ of a granular system subjected to a vertical sinusoidal shaking (with frequency $f$ and amplitude $A$) in a double plate geometry under a confining pressure $p$. Remarkably, similarly to the effective viscosity of our system, the frequency dependence of the friction coefficient $\mu(f)$ reported in this study exhibits a non-monotonic trend. When considering the full $\mu(A,f,p)$ dependence, the authors showed that the friction weakening behavior can be predicted by considering the region of control parameters identified by $\Gamma>1$ and $A^{2}/p>a$ where $a$ is a threshold that depends on the restitution coefficient in the grain-grain interaction. However, the provided criterion only holds in the frequency domain where $\mu(f)$ decreases, thus not including in the analysis its non-monotonic behavior. We want to stress here that our study marks a step forward with respect to Ref. [Clark2023] for two key points. First, when the frequency dependence is not taken into account, we showed that the variable $A^{2}/p$ proposed for the friction-weakening behavior also controls the effective viscosity in our more realistic setup. Moreover, we have also proposed a minimal model for such a scaling in Appendix C. Second, when the full frequency dependence is taken into account, we showed that it is not possible to predict the viscosity reduction by simply considering some thresholds on $A^{2}/p$ and $\Gamma$ because it depends on the control parameters through $K(A,f)$, which is the key variable embodying the non-monotonic dependence on $f$. We believe that the same idea could be used to generalize the analysis on the friction coefficient done in Ref. [Clark2023] by taking into account the non-monotonic $\mu(f)$ dependence. It is also important to point out that the relationship between granular temperature and rheology of vibrated granular materials was already introduced (although with a less systematic analysis) in our previous study [Plati2021Getting] and later in Ref. [irmer2025granular] by Irmer et al. that showed how the granular temperature controls the local rheology in a double plate geometry. However, we point out that the latter study has been performed in a more idealized numerical setup without a direct experimental counterpart and without any specific mention of non-monotonic behaviors of rheological observables and granular temperature.
@@
Finally, inspired by seminal studies on non-vibrated granular flows [forterre2008flows], as an outlook for future development of this research line we highlight the possibility of studying the combined effect of vibration driving parameters and confining pressure also on the volume fraction of vibrated granular systems.
@@ Appendix A Details on numerical simulations
Our DEM simulations make use of the $pair\textunderscore style$ granular command of the LAMMPS package [Plimpton1995, Plimpton2022, LammpsSiteGran] to implement the Hertz–Mindlin [Zhang2005, LammpsSiteGran] model for grain–grain, grain–lid, and grain–vane contact dynamics. The lid is composed of 1773 granular particles glued together, and the rotating vane consists of 4 $\times$ 10 particles glued together and overlapping by half a radius. Grain–wall interactions are described by the same model, implemented with the $wall/gran/region$ command. The Hertz–Mindlin model accounts for both the elastic and dissipative responses during mutual compression of the interacting objects and incorporates the dynamics of their relative translational and rotational motions (see [LammpsSiteGran] for the specific expressions of tangential and normal forces).
@@
Our simulations are initialized by randomly pouring the grains into the container and then letting the lid fall on them. During the pouring process, we impose a high-amplitude shaking to compact the system more effectively. After pouring, the specific $A$ and $f$ to be probed are set, and the torque on the vane is switched on.
@@ Appendix B The role of energy dissipation
In order to support our interpretation based on energy dissipation for the non-monotonic behaviour of $K$, we studied the behaviour of the average work done by the dissipative forces in our system. Our analysis starts by considering that in a time interval $\Delta t$, the variation of the total internal energy in the system (i.e. kinetic plus potiential) is equal to the work $W_{nc}$ done by the non-conservative forces on the system in that interval $U(t+\Delta t)-U(t)=W_{nc}$. Assuming a stationary state and taking the time average of this equality we obtain $\langle W_{nc}\rangle=0$. In other words, different contributions to the work done by non-conservative forces must balance each other. In the setting used to study the granular temperature of the granular medium (i.e. without the rotating vane), we have the work done by the vibrating container $W_{w}$ , the work done by the confining lid $W_{l}$ and finally the work due to the dissipative components in the grain interactions $W_{d}$ (see Sec. A and [LammpsSiteGran]). Thus, assigning the plus sign to the work done on the system and the minus sign to the work done by the system, we have $W_{nc}=W_{w}+W_{l}-W_{d}$. The average work done by internal dissipative forces can then be expressed as $\langle W_{d}\rangle=\langle W_{w}\rangle+\langle W_{l}\rangle.$
@@
where $\omega_{d}$ and $\omega_{w}$ are, respectively, the power associated to the dissipative interactions and the one associated to the vibrating walls. Practically, from our numerical data, we measured the instantaneous force $\bm{F}^{w}_{i}$ between the walls and each grain $i$ at contact for each simulation step $dt$, along with the displacement $\bm{dS}_{i}$ travelled by the grain in the same $dt$. We point out that, to further simplify this numerical measure, we switched off the tangential interactions between the walls and the grains in the simulations performed for this analysis. Then, using Eq. (1) we obtain the per-particle dissipation power
@@
This procedure is much easier than evaluating the dissipation rate calculating the work done by the dissipative contact forces for each couple of interacting grains. Nevertheless, we also performed some auxiliary simulations in a simplified geometry with only a couple of grains to check the consistency of our approach (not shown).
@@ Appendix C A minimal two-block model for the $\bar{A}^{2}/\bar{p}$ scaling
It extends the model introduced in [Plati2021Getting] to account for the kinetic energy measured in DEM simulations, which shows a non-monotonic dependence on the driving frequency with a maximum at $f^{*}$, an overall increase with driving amplitude, and a shift of the peak frequency toward lower values as dissipation grows.
The model is derived from the equation of motion of a single particle, assuming that the collective dynamics of the granular system can be approximated by the behavior of an individual grain. Since the medium remains dense across all shaking regimes, particle motion at short timescales is confined within cages formed by surrounding grains. The key parameter is the mean collision time $\tau_{c}$, which decreases as the driving frequency increases, leading to more frequent dissipative events.
@@ C.1 Results
In Fig. 6 we analyze the dependence of the minimum friction $\mu_{\min}$ on the driving amplitude $A$, for different confining pressures $F_{P}$, comparing Hertzian and Hookean models. As shown in Fig. 6a, at small amplitudes $\mu_{\min}\simeq 1$, while for $A$ larger than a characteristic value (which increases with $F_{P}$), $\mu_{\min}$ decreases monotonically.
@@ C.2 Analytical interpretation of the different collapse of $\mu_{min}$ between Hookean and Hertzian cases
Figure 6 highlights the different scaling behaviors of the minimum friction $\mu_{\min}$ as a function of $A$ and $F_{P}$ for Hookean and Hertzian interactions. In the Hookean case, the scaling is exact: rescaling $A$ and $F_{P}$ by the same factor yields identical results, i.e., the data are strictly invariant under proportional changes of $A$ and $F_{P}$. In contrast, for the Hertzian case the collapse obtained with $A/F_{P}^{2}$ is only approximate and limited to a certain range of parameters, with deviations emerging outside this regime. The origin of this difference can be clarified by analyzing a simplified version of the model, where the stiffness of the bottom spring is taken in the limit $k_{1}\to\infty$. In this regime, the vertical position $z_{1}$ of the lower block is constrained to follow the bottom plate, i.e. $z_{1}=z_{p}+\ell_{0}$, where $z_{p}$ is the vertical position of the bottom plate and $\ell_{0}$ is the rest length of the lower spring. Thus, the dynamics reduce to those of the top block, whose vertical position $z_{2}$ satisfies the equation
@@
Assuming that the bottom plate oscillates as $z_{p}=A\cos(\omega t)$, and introducing the relative distance between the top block and the bottom plate as $z=z_{2}-z_{p}$, Eq. (7) can be rewritten as
$\begin{split}m\ddot{z}+\gamma\dot{z}+k_{2}\left(|z-\ell_{0}|\right)^{\beta}\,\mathrm{sign}(z-\ell_{0})=N-F_{P}\\ +\,mA\omega^{2}\cos(\omega t)+\gamma A\omega\sin(\omega t).\end{split}$ (8)
Introducing dimensionless variables
@@ C.2.1 Hooke case $\beta=1$
Setting, $\beta=1$ in Eq.(9), after rescaling the variable as $\upsilon=y/F^{\prime}_{P}$, we obtain
@@ C.2.2 Hertz case $\beta=3/2$
In this case, we focus on small oscillations around the steady-state solution, i.e. $y\ll 1$. Expanding Eq. (9) up to second order in $y$, we obtain
@@
We next decompose $y(t)=q(t)+\psi(t)$, where $q(t)=\langle y(t)\rangle$ denotes the average of $y(t)$ over one period $T=2\pi/\omega$, and $\psi(t)$ represents the fast fluctuating component. Substituting into Eq. (11) and averaging over one period yields
@@
The evolution equation for $\psi(t)$ can be obtained by subtracting Eq. (12) from Eq. (11). Following the so-called inertial approximation [Ble00], we approximate it as
@@
Substituting this result into Eq. (12), we obtain

### v2

@@ III A unified Picture
Now we point out that the observation of a non-monotonic behavior of $\eta$ as a function of the shaking frequency (Sec. II.1) is in striking contrast with the idea that the rheological response of the granular medium is mainly controlled by $A^{2}/p$. While the viscosity weakening branch can be understood by combining the acceleration threshold $\Gamma=1$ and the $A^{2}/p$ scaling, the raising branch $\eta(\Gamma\geq\Gamma_{\text{m}})$ is not. Indeed, simulations performed at $\Gamma\geq\Gamma_{\text{m}}$ (varied through $f$), with the same amplitude and pressure can manifest effective viscosities that vary over different orders of magnitude, see Fig. 4c. At the same time, we verified that the system’s kinetic energy is only marginally influenced by $p$ (see Appendix A), thus the effect of $p$ on the effective viscosity (see Fig. 4a) can’t be explained with the dependence of $K$ on $p$. Based on our numerical analysis, we define an energy scale related to the confining pressure $K_{p}=p\pi r^{3}$ (i.e. the work necessary to move of a distance $r$ against a constant force $p\pi r^{2}$) and propose the energy ratio $\bar{K}=K/K_{p}$ as the main variable controlling the effective viscosity of the system. In Fig. 4d, we plot $\eta(\Gamma\geq\Gamma_{\text{m}})$ obtained for all the compatible frequencies, amplitudes and pressures varied in our analysis, showing a good collapse on the master curve $\eta\propto\bar{K}^{\alpha}$. Also in this case, our data are well described by a power law $\propto\bar{K}^{\alpha}$, with $\alpha\in[-2,-1.5]$.
@@
We point out that the ratio of energies defining $\bar{K}$ reminds of the competition between thermal agitation and activation energy barriers that characterizes the viscosity divergence in glass-forming supercooled liquids [12] with, however, the important difference that in these systems viscosity divergences are usually exponential or even steeper while in our case we observe a power law. This qualitative analogy with thermal systems is nevertheless conceptually useful to understand the relationship between rheological response and $\bar{K}(A,f,p)$ in our granular setup. A peculiarity of thermal systems is that temperature can be directly controlled by equilibrating the system with a thermal bath. Then, as a consequence of thermodynamic equilibrium, also the average kinetic energy of the system is fixed and proportional to the temperature. Instead, in a granular system, the average kinetic energy $K(A,f)$ depends non-trivially on the driving parameters. In other words, the relationship between injected and adsorbed energy is much more complex when thermodynamic equilibrium is violated, as in granular systems. Therefore, to understand how the driving parameters $A$ and $f$ control the rheology of a granular system it is important to first clarify how they act on its granular temperature $K$. Our numerical analysis shows that, since $\eta$ depends non-monotonously on $f$, there is no simple scaling in the form $\propto A^{\beta}f^{\lambda}$, with $\beta$ and $\lambda$ arbitrary exponents, that could predict the behavior of the effective viscosity. Panel c and d of Fig. 4 show in fact that $\eta$ depends on $A$ and $f$ through $K(A,f)$, which embodies the non-monotonic dependence on $f$.
@@ IV Conclusions and link with previous studies
In this paper we investigated the rheology of a dense vibrated granular system by means of realistic numerical DEM simulations of a stress imposed vane rheometry setup. We varied the vibration amplitude $A$ and frequency $f$ of the vertical sinusoidal shaking as well as the confining pressure $p$. We found that the variable controlling the effective viscosity is $\bar{K}=K/K_{p}$ where $K(A,f)$ is the typical agitation energy of a grain in the cage formed by the surrounding particles and $K_{p}\propto p$ is an energy scale set by the confining pressure that could be interpreted as the energy needed to escape from such a cage. Our proposed scaling takes into account the non-monotonic effect of $f$ on $\eta$ through the energy transfer mechanism discussed in Ref. [29] and Appendix B.
@@
We now discuss the connection between our study and previous works on the rheology of vibrated granular materials in simplified geometries. We start by referring to the numerical study reported in [7] which focuses on the macroscopic friction coefficient $\mu$ of a granular system subjected to a vertical sinusoidal shaking (with frequency $f$ and amplitude $A$) in a double plate geometry under a confining pressure $p$. Remarkably, similarly to the effective viscosity of our system, the frequency dependence of the friction coefficient $\mu(f)$ reported in this study exhibits a non-monotonic trend. When considering the full $\mu(A,f,p)$ dependence, the authors showed that the friction weakening behavior can be predicted by considering the region of control parameters identified by $\Gamma>1$ and $A^{2}/p>a$ where $a$ is a threshold that depends on the restitution coefficient in the grain-grain interaction. However, the provided criterion only holds in the frequency domain where $\mu(f)$ decreases, thus not including in the analysis its non-monotonic behavior. We want to stress here that our study marks a step forward with respect to Ref. [7] for two key points. First, when the frequency dependence is not taken into account, we showed that the variable $A^{2}/p$ proposed for the friction-weakening behavior also controls the effective viscosity in our more realistic setup. Moreover, we have also proposed a minimal model for such a scaling in Appendix C. Second, when the full frequency dependence is taken into account, we showed that it is not possible to predict the viscosity reduction by simply considering some thresholds on $A^{2}/p$ and $\Gamma$ because it depends on the control parameters through $K(A,f)$, which is the key variable embodying the non-monotonic dependence on $f$. We believe that the same idea could be used to generalize the analysis on the friction coefficient done in Ref. [7] by taking into account the non-monotonic $\mu(f)$ dependence. It is also important to point out that the relationship between granular temperature and rheology of vibrated granular materials was already introduced (although with a less systematic analysis) in our previous study [29] and later in Ref. [23] by Irmer et al. that showed how the granular temperature controls the local rheology in a double plate geometry. However, we point out that the latter study has been performed in a more idealized numerical setup without a direct experimental counterpart and without any specific mention of non-monotonic behaviors of rheological observables and granular temperature.
@@
For future studies, it would be useful to develop a theoretical framework to estimate the exponent $\alpha$ of the power-law that links $\eta$ and $\bar{K}$. Moreover, inspired by seminal studies on non-vibrated granular flows [17], it would be interesting to study the combined effect of vibration driving parameters and confining pressure also on the volume fraction of vibrated granular systems.
@@ Acknowledgements.
Data Availability
The data that support the findings of this article are openly available [2].
@@ Appendix A Details on numerical simulations
Our DEM simulations make use of the $pair\textunderscore style$ granular command of the LAMMPS package [30, 32, 1] to implement the Hertz–Mindlin [35, 1] model for grain–grain, grain–lid, and grain–vane contact dynamics. The lid is composed of 1773 granular particles glued together, and the rotating vane consists of 4 $\times$ 10 particles glued together and overlapping by half a radius. Grain–wall interactions are described by the same model, implemented with the $wall/gran/region$ command. The Hertz–Mindlin model accounts for both the elastic and dissipative responses during mutual compression of the interacting objects and incorporates the dynamics of their relative translational and rotational motions (see [1] for the specific expressions of tangential and normal forces).
@@
Our simulations are initialized by randomly pouring the grains into the container and then by letting the lid fall on them. During the pouring process, we impose a high-amplitude shaking to compact the system more effectively. After pouring, the specific $A$ and $f$ to be probed are set, and the torque on the vane is switched on.
[…]
Figure 5: Kinetic energy of the granular medium as a function of $p$ for four shaking amplitudes and $f=400$ Hz.
@@ Appendix B The role of energy dissipation
In order to support our interpretation based on energy dissipation for the non-monotonic behaviour of $K$, we studied the behaviour of the average work done by the dissipative forces in our system. Our analysis starts by considering that in a time interval $\Delta t$, the variation of the total internal energy in the system (i.e. kinetic plus potiential) is equal to the work $W_{nc}$ done by the non-conservative forces on the system in that interval $U(t+\Delta t)-U(t)=W_{nc}$. Assuming a stationary state and taking the time average of this equality we obtain $\langle W_{nc}\rangle=0$. In other words, different contributions to the work done by non-conservative forces must balance each other. In the setting used to study the granular temperature of the granular medium (i.e. without the rotating vane), we have the work done by the vibrating container $W_{w}$ , the work done by the confining lid $W_{l}$ and finally the work due to the dissipative components in the grain interactions $W_{d}$ (see Sec. A and [1]). Thus, assigning the plus sign to the work done on the system and the minus sign to the work done by the system, we have $W_{nc}=W_{w}+W_{l}-W_{d}$. The average work done by internal dissipative forces can then be expressed as $\langle W_{d}\rangle=\langle W_{w}\rangle+\langle W_{l}\rangle.$
@@
[…]
@@
This procedure is much easier than evaluating the dissipation rate from the work done by the dissipative contact forces for each couple of interacting grains. Nevertheless, we also performed some auxiliary simulations in a simplified geometry with only a couple of grains to check the consistency of our approach (not shown).
@@ Appendix C A minimal two-block model for the $\bar{A}^{2}/\bar{p}$ scaling
It extends the model introduced in [29] to account for the kinetic energy measured in DEM simulations, which shows a non-monotonic dependence on the driving frequency with a maximum at $f^{*}$, an overall increase with driving amplitude, and a shift of the peak frequency toward lower values as dissipation grows.
The model is derived from the equation of motion of a single particle, by assuming that the collective dynamics of the granular system can be approximated by the behavior of an individual grain. Since the medium remains dense across all shaking regimes, particle motion at short timescales is confined within cages formed by surrounding grains. The key parameter is the mean collision time $\tau_{c}$, which decreases as the driving frequency increases, leading to more frequent dissipative events.
@@ C.1 Results
In Fig. 8 we analyze the dependence of the minimum friction $\mu_{\min}$ on the driving amplitude $A$, for different confining pressures $F_{P}$, by comparing Hertzian and Hookean models. As shown in Fig. 8a, at small amplitudes $\mu_{\min}\simeq 1$, while for $A$ larger than a characteristic value (which increases with $F_{P}$), $\mu_{\min}$ decreases monotonically.
@@ C.2 Analytical interpretation of the different collapse of $\mu_{min}$ between Hookean and Hertzian cases
Figure 8 highlights the different scaling behaviors of the minimum friction $\mu_{\min}$ as a function of $A$ and $F_{P}$ for Hookean and Hertzian interactions. In the Hookean case, the scaling is exact: by rescaling $A$ and $F_{P}$ by the same factor yields identical results, i.e., the data are strictly invariant under proportional changes of $A$ and $F_{P}$. In contrast, for the Hertzian case the collapse obtained with $A/F_{P}^{2}$ is only approximate and limited to a certain range of parameters, with deviations emerging outside this regime. The origin of this difference can be clarified by analyzing a simplified version of the model, where the stiffness of the bottom spring is taken in the limit $k_{1}\to\infty$. In this regime, the vertical position $z_{1}$ of the lower block is constrained to follow the bottom plate, i.e. $z_{1}=z_{p}+\ell_{0}$, where $z_{p}$ is the vertical position of the bottom plate and $\ell_{0}$ is the rest length of the lower spring. Thus, the dynamics reduce to those of the top block, whose vertical position $z_{2}$ satisfies the equation
@@
By assuming that the bottom plate oscillates as $z_{p}=A\cos(\omega t)$, and by introducing the relative distance between the top block and the bottom plate as $z=z_{2}-z_{p}$, Eq. (7) can be rewritten as
$split m\ddot{z}+\gamma\dot{z}+k_{2}\left(|z-\ell_{0}|\right)^{\beta}\,\mathrm{sign}(z-\ell_{0})=N-F_{P}\\ +\,mA\omega^{2}\cos(\omega t)+\gamma A\omega\sin(\omega t).$ (8)
By introducing dimensionless variables
@@ C.2.1 Hooke case $\beta=1$
By setting, $\beta=1$ in Eq.(9), after rescaling the variable as $\upsilon=y/F^{\prime}_{P}$, we obtain
@@ C.2.2 Hertz case $\beta=3/2$
In this case, we focus on small oscillations around the steady-state solution, i.e. $y\ll 1$. By expanding Eq. (9) up to second order in $y$, we obtain
@@
We next decompose $y(t)=q(t)+\psi(t)$, where $q(t)=\langle y(t)\rangle$ denotes the average of $y(t)$ over one period $T=2\pi/\omega$, and $\psi(t)$ represents the fast fluctuating component. By substituting $y(t)$ into Eq. (11) and averaging over one period yields
@@
The evolution equation for $\psi(t)$ can be obtained by subtracting Eq. (12) from Eq. (11). By following the so-called inertial approximation [5], we approximate it as
@@
By substituting this result into Eq. (12), we obtain